The value of cos 22.5 is.........????????
 
Plxxx answer fast

Grade:11

2 Answers

Arun
4 years ago
 
Dear Student,
1+cos 45^{\circ}= 2 cos^222.5^{\circ}    using 1 + cos 2A = 2 cos^2A
so    cos^222.5^{\circ}= (\sqrt{2}+ 1)/2\sqrt{2}
so   cos22.5^{\circ}= \sqrt{(\sqrt{2}+ 1)/2\sqrt{2}}     
 
 
Hope it helps
In case of any difficulty, please feel free to ask gain.
 
Regards
Arun (askIITians forum expert)
sushmitha raj
4 years ago
cosA=2cos2A/2-1
hereA=45
cos45=2cos222.5-1
cos22.5=(21/2+1/81/2)½
hope it helps
regards 
all the best for your IIT exam